# Zhuo Zhang College of Computer Science and Technology

### v2

@@ 5.2.3 Conformal Prediction for Distribution-Free UQ
Recent breakthrough: conformal prediction provides coverage guarantees without distributional assumptions (Staber2024).
@@ 5.2.4 Uncertainty Propagation Through Parametric PDEs
Multi-fidelity Monte Carlo achieves optimal bias-variance tradeoff (Peherstorfer2018).
@@ 5.3 Meta-Learning and Rapid Adaptation
Meta-learning (Finn2017; Yin2022; Huang2022) enables neural methods to quickly adapt to new parameter regimes or PDE types with minimal retraining—crucial when parameter distributions shift or new physics are encountered.
@@ 5.3.1 Model-Agnostic Meta-Learning (MAML) for PDEs
MAML (Finn2017) learns initialization $\theta_{0}$ such that few gradient steps on new tasks yield good performance.
@@ 5.3.2 Transfer Learning Across Parameter Regimes
Results: Desai et al. (Desai2021) demonstrated: Transfer from $\text{Re}=100$ to $\text{Re}=1000$: 80% reduction in training iterations, Maintains 95% of full-training accuracy, and Critical: transferred representations capture flow physics invariant to Re
@@ 5.3.3 Meta-Auto-Decoder for Parametric PDEs
Yin et al. (Yin2022) developed Meta-Auto-Decoder combining meta-learning with latent representations:
@@ 5.3.4 Continual Learning for Evolving Parameter Distributions
Application - Climate Modeling: As climate changes, parameter distributions (e.g., atmospheric CO2, temperature patterns) evolve. Continual learning enables neural operators to update without forgetting historical patterns (Chattopadhyay2023).
@@ 5.4.2 Hybrid Physics-ML Models (PINO Framework)
PINO (Li2024PINO) combines data-driven learning with physics-informed losses at multiple resolutions:
@@
PINO achieves best of both worlds: data efficiency from physics, accuracy from multi-fidelity (Penwarden2023; Howard2023; Peherstorfer2018) data.
@@ 5.5.1 The Foundation Model Paradigm
5.5.2 Representative Approaches
1. Poseidon: Hao et al. (Hao2023Poseidon) pre-trained on 15 PDE families: Architecture: Transformer-based operator with 100M parameters, Pre-training: 1M PDE solutions across equations, parameters, domains, Fine-tuning: 10-100 samples for new PDEs, and Performance: Zero-shot transfer to new PDEs with $10\%$ error; fine-tuning achieves $3\%$
2. DPOT (Deep Physical Operator Transformer): Another foundation model focusing on physical priors: Embeds conservation laws in architecture, Trained on synthetic + real physics data, and Generalizes across spatial dimensions (1D, 2D, 3D)
3. MOFS (Multi-Operator Foundation Solver): Specializes in operator composition: Learns elemental operators (advection, diffusion, reactions), Composes for complex PDEs, and Analogous to “building blocks” in language models
@@ 6.1.3 Neuraloperator
Strengths: Cutting-edge architectures, research-oriented, excellent for prototyping Limitations: Less documentation than DeepXDE (Lu2021), fewer built-in examples
@@ 6.1.4 Comparative Overview
DeepXDE (Lu2021) PINN, DeepONet High Medium Research, education, prototyping
@@ 6.2.1 PDEBench
Reference: Takamoto et al. (Takamoto2022) Repository: github.com/pdebench/PDEBench
@@ 1. Incomplete Convergence Guarantees
Evidence: Krishnapriyan et al. (Krishnapriyan2021) systematically documented failure modes: Diffusion-dominated vs. advection-dominated regime switching causes training collapse, Multi-scale problems exhibit oscillatory loss with no convergence, and Spectral bias prevents learning high-frequency components
@@ 1. Training Instability and Hyperparameter Sensitivity
Current Approaches: Learning rate annealing, gradient balancing (Wang2021Grad), uncertainty weighting (Kendall2018)—but no universal solution.
@@ 7.2.2 Direction 2: Geometry-Parameterized Problems
Target Applications: Automotive design: Optimize aerodynamics across full design space, Medical imaging: Patient-specific simulations from scan to diagnosis in minutes, and Zhu et al. (Zhu2023PhaseField) introduced Phase-Field DeepONet using energy-based loss functions for pattern formation, enabling fast Allen-Cahn and Cahn-Hilliard simulations. Lee et al. (Lee2025FEONO) developed FE Operator Networks for high-dimensional elasticity (d=50 parameters), achieving 60% error reduction.
@@ 7.2.4 Direction 4: Uncertainty Quantification and Reliability
Promising Approaches: Conformal prediction: Distribution-free coverage guarantees (breakthrough: Staber et al. 2024)—extend to sequential predictions and multi-fidelity (Penwarden2023; Howard2023; Peherstorfer2018) settings., Bayesian neural operators: Make tractable via variational inference, Laplace approximation, or ensemble distillation., Certified robustness: Borrow techniques from adversarial ML—prove that small parameter perturbations cause bounded solution changes., and Physics-based validation: Check predictions against conservation laws, positivity constraints, maximum principles as sanity checks.
@@ 3. Equation Discovery:
Learn governing equations from data using neural operators as differentiable simulators (Raissi2018; Brunton2016; Zhang2022; Geneva2020; Chen2019NeuralODE): Sparse regression (SINDy) (Brunton2016) with neural-operator-generated training data, Neural differential equations (Chen2019NeuralODE) with interpretable coefficients, and Symbolic regression guided by neural operator gradients (Zhang2022)
@@ 1. Standardized Benchmarks
Establish comprehensive benchmark suite (Takamoto2022) covering: Diverse PDE types (elliptic, parabolic, hyperbolic, mixed), Parameter dimensions from 1 to 100+, Multiple application domains, and Varying difficulty levels
@@ 7.4 Context and Related Surveys
This survey builds upon and extends several complementary surveys and foundational works in the field. Comprehensive overviews of physics-informed machine learning have been provided by Cuomo et al. (Cuomo2022), Azizzadenesheli et al. (Azizzadenesheli2024), and Karniadakis et al., establishing the broader context for physics-informed neural networks. Theoretical foundations in deep learning for PDEs were laid by Han et al. (Han2018), E and Yu (Weinan2019), Sirignano and Spiliopoulos (Sirignano2018), and Berg and Nyström (Berg2018), demonstrating feasibility of neural approximations for high-dimensional problems.
Reduced-order modeling foundations relevant to parametric PDEs are comprehensively covered by Hesthaven et al. (Hesthaven2016), Berkooz et al. (Berkooz1993), and Peherstorfer et al. (Peherstorfer2016), providing mathematical context for solution manifold structure. Methodological comparisons and benchmarking efforts (Lu2022; Geneva2022; Gao2021; Fuks2020) have critically evaluated trade-offs between different neural approaches.
Domain-specific advances complement our parametric focus: geophysical applications (Bihlo2022), quantum mechanics (Hermann2020), chemical kinetics (Ji2021), manufacturing (Zobeiry2021; Cai2022), nanophotonics (Wiecha2021), and financial mathematics (Ruf2020) demonstrate breadth of applicability. Meta-learning and adaptive approaches (Huang2022; Huang2023; Hao2023; Geneva2020; Goswami2020) address rapid adaptation challenges critical for parametric problems.
Advanced architectural innovations including graph-based methods (Hao2023GNOT2; Gao2021), Hamiltonian-preserving networks (Greydanus2019), wavelet-based operators (Gupta2021), factorized representations (Tran2023), and geometric approaches (Baque2018) continue expanding the toolkit. Multi-fidelity strategies (Howard2023; Penwarden2023) and enhanced conservation law enforcement (Jagtap2021) address data efficiency and physical consistency.
@@ 8 Conclusion
This comprehensive survey has examined the rapid evolution of neural methods for solving parametric partial differential equations, with emphasis on physics-informed neural networks and neural operators. The field has matured remarkably since the seminal works of Raissi et al. (Raissi2019) and Lu et al. (Lu2021), transitioning from proof-of-concept demonstrations to practical engineering tools achieving $10^{3}$-$10^{5}\times$ computational speedups while maintaining or exceeding traditional solver accuracy.
@@ 8.1.2 Theoretical Understanding
Universal Approximation: Rigorous proofs establish that neural operators can approximate any continuous operator between function spaces (Kovachki2023; Lu2021), providing mathematical legitimacy to the approach.
@@
Failure Mode Analysis: Systematic studies (Krishnapriyan2021; Wang2022NTK) have identified when and why physics-informed methods struggle (spectral bias, stiff equations, multi-scale problems), guiding algorithm development and honest assessment of applicability.
Uncertainty Quantification: The advent of conformal prediction for PDEs (Staber2024) provides distribution-free coverage guarantees, addressing a critical need for reliable uncertainty estimates in safety-critical applications.
@@ 8.1.3 Cross-Domain Impact
Solid Mechanics: Zhu et al. (Zhu2023PhaseField) introduced Phase-Field DeepONet using energy-based loss functions for pattern formation, enabling fast Allen-Cahn and Cahn-Hilliard simulations. Lee et al. (Lee2025FEONO) developed FE Operator Networks for high-dimensional elasticity (d=50 parameters), achieving 60% error reduction.
@@
Electromagnetics: Metamaterial inverse design (Lu2021Adaptive; Meng2022), antenna optimization, and Maxwell equation solving for varying material properties demonstrate the technology’s versatility across physics domains.
@@ 8.3.2 For Practitioners
1. Start with Established Methods: For production applications, prioritize mature frameworks (DeepXDE (Lu2021), NVIDIA Modulus) with community support and track records. Avoid cutting-edge methods until well-validated.

### v3

@@ 5.2.3 Conformal Prediction for Distribution-Free UQ
Recent breakthrough: conformal prediction provides coverage guarantees without distributional assumptions (Chalot et al., 2023).
@@ 5.2.4 Uncertainty Propagation Through Parametric PDEs
Multi-fidelity Monte Carlo achieves optimal bias-variance tradeoff (Peherstorfer et al., 2018).
@@ 5.3 Meta-Learning and Rapid Adaptation
Meta-learning (Finn et al., 2017; Yin et al., 2022; Ye et al., 2024) enables neural methods to quickly adapt to new parameter regimes or PDE types with minimal retraining—crucial when parameter distributions shift or new physics are encountered.
@@ 5.3.1 Model-Agnostic Meta-Learning (MAML) for PDEs
MAML (Finn et al., 2017) learns initialization $\theta_{0}$ such that few gradient steps on new tasks yield good performance.
@@ 5.3.2 Transfer Learning Across Parameter Regimes
Results: Desai et al. (Desai et al., 2021) demonstrated: Transfer from $\text{Re}=100$ to $\text{Re}=1000$: 80% reduction in training iterations, Maintains 95% of full-training accuracy, and Critical: transferred representations capture flow physics invariant to Re
@@ 5.3.3 Meta-Auto-Decoder for Parametric PDEs
Yin et al. (Yin et al., 2022) developed Meta-Auto-Decoder combining meta-learning with latent representations:
@@ 5.3.4 Continual Learning for Evolving Parameter Distributions
Application - Climate Modeling: As climate changes, parameter distributions (e.g., atmospheric CO2, temperature patterns) evolve. Continual learning enables neural operators to update without forgetting historical patterns (Chattopadhyay et al., 2022).
@@ 5.4.2 Hybrid Physics-ML Models (PINO Framework)
PINO (Li et al., 2024b) combines data-driven learning with physics-informed losses at multiple resolutions:
@@
PINO achieves best of both worlds: data efficiency from physics, accuracy from multi-fidelity (Penwarden et al., 2022; Lu et al., 2022b; Peherstorfer et al., 2018) data.
@@ 6.1.3 Neuraloperator
Strengths: Cutting-edge architectures, research-oriented, excellent for prototyping Limitations: Less documentation than DeepXDE (Lu et al., 2021a), fewer built-in examples
@@ 6.1.4 Comparative Overview
DeepXDE (Lu et al., 2021a) PINN, DeepONet High Medium Research, education, prototyping
@@ 6.2.1 PDEBench
Reference: Takamoto et al. (Takamoto et al., 2022) Repository: github.com/pdebench/PDEBench
@@ 1. Incomplete Convergence Guarantees
Evidence: Krishnapriyan et al. (Krishnapriyan et al., 2021) systematically documented failure modes: Diffusion-dominated vs. advection-dominated regime switching causes training collapse, Multi-scale problems exhibit oscillatory loss with no convergence, and Spectral bias prevents learning high-frequency components
@@ 1. Training Instability and Hyperparameter Sensitivity
Current Approaches: Learning rate annealing, gradient balancing (Schäfer and Anandkumar, 2019), uncertainty weighting (Kendall et al., 2018)—but no universal solution.
@@ 7.2.2 Direction 2: Geometry-Parameterized Problems
Target Applications: Automotive design: Optimize aerodynamics across full design space, Medical imaging: Patient-specific simulations from scan to diagnosis in minutes, and Zhu et al. (Li et al., 2023a) introduced Phase-Field DeepONet using energy-based loss functions for pattern formation, enabling fast Allen-Cahn and Cahn-Hilliard simulations. Lee et al. (Lee et al., 2025) developed FE Operator Networks for high-dimensional elasticity (d=50 parameters), achieving 60% error reduction.
@@ 7.2.4 Direction 4: Uncertainty Quantification and Reliability
Promising Approaches: Conformal prediction: Distribution-free coverage guarantees (breakthrough: Staber et al. 2024)—extend to sequential predictions and multi-fidelity (Penwarden et al., 2022; Lu et al., 2022b; Peherstorfer et al., 2018) settings., Bayesian neural operators: Make tractable via variational inference, Laplace approximation, or ensemble distillation., Certified robustness: Borrow techniques from adversarial ML—prove that small parameter perturbations cause bounded solution changes., and Physics-based validation: Check predictions against conservation laws, positivity constraints, maximum principles as sanity checks.
@@ 3. Equation Discovery:
Learn governing equations from data using neural operators as differentiable simulators (Raissi, 2018; Brunton et al., 2016; Zhang and Lin, 2018; Geneva and Zabaras, 2020; Chen et al., 2018): Sparse regression (SINDy) (Brunton et al., 2016) with neural-operator-generated training data, Neural differential equations (Chen et al., 2018) with interpretable coefficients, and Symbolic regression guided by neural operator gradients (Zhang and Lin, 2018)
@@ 1. Standardized Benchmarks
Establish comprehensive benchmark suite (Takamoto et al., 2022) covering: Diverse PDE types (elliptic, parabolic, hyperbolic, mixed), Parameter dimensions from 1 to 100+, Multiple application domains, and Varying difficulty levels
@@ 7.4 Context and Related Surveys
This survey builds upon and extends several complementary surveys and foundational works in the field. Comprehensive overviews of physics-informed machine learning have been provided by Cuomo et al. (Cuomo et al., 2022), Azizzadenesheli et al. (Azizzadenesheli et al., 2024), and Karniadakis et al., establishing the broader context for physics-informed neural networks. Theoretical foundations in deep learning for PDEs were laid by Han et al. (Han et al., 2018), Yu (Yu and others, 2018), Sirignano(Sirignano and Spiliopoulos, 2018), and Berg(Berg and Nyström, 2018), demonstrating feasibility of neural approximations for high-dimensional problems.
Reduced-order modeling foundations relevant to parametric PDEs are comprehensively covered by Hesthaven et al. (Hesthaven et al., 2016), Berkooz et al. (Berkooz et al., 1993), and Peherstorfer et al. (Peherstorfer et al., 2014), providing mathematical context for solution manifold structure. Methodological comparisons and benchmarking efforts (Lu et al., 2022a; Geneva and Zabaras, 2022; Gao et al., 2022; Fuks and Tchelepi, 2020) have critically evaluated trade-offs between different neural approaches.
Domain-specific advances complement our parametric focus: geophysical applications (Bihlo and Popovych, 2022), quantum mechanics (Hermann et al., 2020), chemical kinetics (Ji et al., 2021), manufacturing (Zobeiry and Humfeld, 2021; Cai et al., 2022), nanophotonics (Wiecha and Muskens, 2020), and financial mathematics (Ruf and Wang, 2020) demonstrate breadth of applicability. Meta-learning and adaptive approaches (Ye et al., 2024; Chiniforooshan Esfahani, 2023; Hao et al., 2022; Geneva and Zabaras, 2020; Goswami et al., 2020a) address rapid adaptation challenges critical for parametric problems.
Advanced architectural innovations including graph-based methods (Hao et al., 2023; Gao et al., 2022), Hamiltonian-preserving networks (Greydanus et al., 2019), wavelet-based operators (Gupta et al., 2021), factorized representations (Tran et al., 2021), and geometric approaches (Masci et al., 2015) continue expanding the toolkit. Multi-fidelity strategies (Lu et al., 2022b; Penwarden et al., 2022) and enhanced conservation law enforcement (Jagtap et al., 2020) address data efficiency and physical consistency.
@@ 8 Conclusion
This comprehensive survey has examined the rapid evolution of neural methods for solving parametric partial differential equations, with emphasis on physics-informed neural networks and neural operators. The field has matured remarkably since the seminal works of Raissi et al. (Raissi et al., 2019) and Lu et al. (Lu et al., 2021a), transitioning from proof-of-concept demonstrations to practical engineering tools achieving $10^{3}$-$10^{5}\times$ computational speedups while maintaining or exceeding traditional solver accuracy.
@@ 8.1.2 Theoretical Understanding
Universal Approximation: Rigorous proofs establish that neural operators can approximate any continuous operator between function spaces (Kovachki et al., 2023a; Lu et al., 2021a), providing mathematical legitimacy to the approach.
@@
Failure Mode Analysis: Systematic studies (Krishnapriyan et al., 2021; Wang et al., 2022b) have identified when and why physics-informed methods struggle (spectral bias, stiff equations, multi-scale problems), guiding algorithm development and honest assessment of applicability.
Uncertainty Quantification: The advent of conformal prediction for PDEs (Chalot et al., 2023) provides distribution-free coverage guarantees, addressing a critical need for reliable uncertainty estimates in safety-critical applications.
@@ 8.1.3 Cross-Domain Impact
Solid Mechanics: Zhu et al. (Li et al., 2023a) introduced Phase-Field DeepONet using energy-based loss functions for pattern formation, enabling fast Allen-Cahn and Cahn-Hilliard simulations. Lee et al. (Lee et al., 2025) developed FE Operator Networks for high-dimensional elasticity (d=50 parameters), achieving 60% error reduction.
@@
Electromagnetics: Metamaterial inverse design (Lu et al., 2021b; Meng et al., 2022), antenna optimization, and Maxwell equation solving for varying material properties demonstrate the technology’s versatility across physics domains.
@@ 8.3.2 For Practitioners
1. Start with Established Methods: For production applications, prioritize mature frameworks (DeepXDE (Lu et al., 2021a), NVIDIA Modulus) with community support and track records. Avoid cutting-edge methods until well-validated.